OpenStudy (anonymous):

Simplification of expressions. (a^2 over b^3)^4

OpenStudy (anonymous):

32/ab^3

OpenStudy (anonymous):

\[a{4}/b{6}\]

OpenStudy (mysesshou):

\[(\frac{a^2}{b^3})^4\] = (a^2/b^3)^4 a power to a power, multiplies, giving (a^8)/(b^12)

OpenStudy (anonymous):

sorry I misread the power. thought it everything to the power of 2. mysesshou clears that out